### v1

@@ I Introduction
In most optical magnetometry experiments, the polarization across the light beam is approximately uniform. Meanwhile, recent advances in optics made it possible to generate light fields with space-varying polarization Rubinsztein-Dunlop et al. (2017). The best known examples of such vector beams include radially and azimuthally polarized beams with an azimuthally varying linear polarization surrounding an optical vortex Gbur (2017). These spatially varying light polarizations can excite locally varying magnetization profiles in atoms Wang et al. (2020). It has been demonstrated recently in cold Castellucci et al. (2021) and warm Qiu et al. (2021); Maldonado et al. (2024) atomic vapors to test the magnetic field components transverse and along the optical axis simultaneously. These experiments determined static magnetic fields from absorption profiles of a vector beam after its passage through the rubidium vapor.
Beyond static magnetic fields, the detection of time-dependent magnetic fields is important, especially in the radio-frequency domain. Detection of fields in the kilohertz to gigahertz frequency range finds many applications, from radio communication to detection of nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) signals Savukov et al. (2005). An example of an atomic magnetometer for detection of radio-frequency magnetic fields is magnetometer Ledbetter et al. (2007) based on a nonlinear magneto-optical rotation. In this paper, we explore the possibility of detecting oscillating (radio-frequency) magnetic fields based on another effect, namely the dependence of the absorption profile of a vector beam, propagating through an atomic vapor, on the strength and frequency of the magnetic field. In addition to the light field and the magnetic field to be measured, here we need an additional reference static magnetic field with well-known direction and strength. As in previous works Savukov et al. (2005); Ledbetter et al. (2007), the static magnetic field is perpendicular to the oscillating one. While, however, the static field in Refs. Savukov et al. (2005); Ledbetter et al. (2007) was added to achieve the Zeeman resonance of atoms, it is here applied in order to trigger oscillations in the orientation of the resulting magnetic field relative to the light polarization.
@@
To determine the effect of the applied magnetic field on the light absorption profile, we employ density matrix theory, whose basic formulas are briefly reviewed in Sec. II. We show, in particular, how to calculate the time evolution of the atomic density matrix from the transition amplitudes for vector beams and Zeeman shifts caused by the superposition of the reference and test magnetic fields. From this density matrix we obtain in Sec. III the light absorption profile illustrated for the case of the 87Rb D2222 line. We first investigate the time evolution of the absorption profiles and show that the petal-like absorption patterns rotate about the beam axis at a rate that depends on the applied AC magnetic field. In general, this effect can be used for magnetometry, but it requires a high time resolution of the detector. An easier characteristic to observe is the absorption profile averaged over the measurement time. We found that even this profile can be sensitive to both the strength and frequency of the test magnetic field, and this sensitivity is most pronounced for frequencies up to a hundred kHz and strengths up to several Gauss for the reference field of about one Gauss. Finally, Sec. IV provides a brief summary and outlook.
@@ II.2 Vector light beams
where $A_{0}=-E_{0}\omega$ and $a_{\varkappa m_{\gamma}}(\bm{k}_{\bot})$ is a weight function given by:
@@ II.3 Transition amplitudes
with $\bm{F}=\bm{I}+\bm{J}$, where $\bm{I}$ and $\bm{J}$ are the nuclear and electron angular momenta, respectively, $M$ is the projection of $\bm{F}$ on the quantization axis, and $\alpha$ denotes all additional quantum numbers required to specify the state uniquely. Moreover, $q$ runs over all electrons in a target atom and $\bm{\alpha}_{q}$ denotes the vector of Dirac matrices for the $q$th particle Johnson (2007). In Eq. (8), we have introduced the impact parameter $\bm{b}=(b\cos\phi_{b},b\sin\phi_{b},0)$ to specify the position of the atom within the wave front Knyazev and Serbo (2018). In particular, $b=0$ corresponds to an atom located on the vortex line of a light beam. The impact parameter plays an important role in the analysis due to the complex spatial structure of the vector beam, see Fig. 3.
@@
where we have used the notation $\langle\alpha_{e}J_{e}||H_{\gamma}(pL)||\alpha_{g}J_{g}\rangle$ to denote the reduced matrix element for magnetic ($p=0$) and electric ($p=1$) transitions. Furthermore, the Euler angles as the arguments of the Wigner $D$-functions $d^{L}_{M,\lambda}(\theta_{k})$ and $D^{L}_{M_{e}-M_{g},M}(\pi,\pi/2,\pi-\phi_{B}(t))$ in Eq. (10) characterize the rotation from the atomic frame with the quantization axis along the magnetic field $\bm{B}_{\mathrm{total}}(t)$ to the photon frame with the quantization axis along the wave vector $\bm{k}$ Rose (1957). The time-dependent angle $\pi-\phi_{B}(t)$ refers to the oscillation of the polarization vector of the incident light at the position of the atom in its (time-dependent) reference frame.

### v2

@@ I Introduction
In most optical magnetometry experiments, the polarization across the light beam is approximately uniform. Meanwhile, recent advances in optics made it possible to generate light fields with space-varying polarization Rubinsztein-Dunlop et al. (2017). The best known examples of such vector beams include radially and azimuthally polarized beams with an azimuthally varying linear polarization surrounding an optical vortex Gbur (2017). These spatially varying light polarizations can excite locally varying magnetization profiles in atoms Wang et al. (2020). It has been demonstrated recently in cold Castellucci et al. (2021) and warm Qiu et al. (2021); Maldonado et al. (2024) atomic vapors that such spatially varying light-matter interaction allows a simultaneous measurement of magnetic field components transverse and along the optical axis. These experiments determined static magnetic fields from absorption profiles of a vector beam after its passage through the rubidium vapor. We note that in this work as well as for the mechanisms proposed here the properties of the vector light change over length scales much larger than the extend of individual atoms.
Beyond static magnetic fields, the detection of time-dependent magnetic fields is important, especially in the radio-frequency domain. Detection of fields in the kilohertz to gigahertz frequency range finds many applications, from radio communication to detection of nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) signals Savukov et al. (2005). An example of an atomic magnetometer for detection of radio-frequency magnetic fields is magnetometer Ledbetter et al. (2007) based on a nonlinear magneto-optical rotation. In this paper, we explore the possibility of detecting oscillating (radio-frequency) magnetic fields based on another effect, namely the dependence of the absorption profile of a vector beam, propagating through an atomic vapor, on the strength and frequency of the magnetic field. In addition to the light field and the magnetic field to be measured, here we need an additional reference static magnetic field with well-known direction and strength. As in previous works Savukov et al. (2005); Ledbetter et al. (2007), the static magnetic field is perpendicular to the oscillating one. While the static field in Refs. Savukov et al. (2005); Ledbetter et al. (2007) was added to achieve the Zeeman resonance of atoms, it is here applied in order to trigger oscillations in the orientation of the resulting magnetic field relative to the light polarization.
@@
To determine the effect of the applied magnetic field on the light absorption profile, we employ density matrix theory, whose basic formulas are briefly reviewed in Sec. II. We show, in particular, how to calculate the time evolution of the atomic density matrix from the transition amplitudes for vector beams and Zeeman shifts caused by the superposition of the reference and test magnetic fields. We assume that the amplitude and polarization of the vector light change over distances much larger than an individual atom, so that we can evaluate light-matter interaction locally. Furthermore, we consider an adiabatic regime, where atomic dynamics happen at time scales much faster than changes to the light intensity (as may occur for pulsed light), and changes to the direction and intensity of the magnetic field. From the density matrix we obtain in Sec. III the light absorption profile illustrated for the case of the 87Rb D2222 line. We first investigate the time evolution of the absorption profiles and show that the petal-like absorption patterns rotate about the beam axis at a rate that depends on the applied AC magnetic field. In general, this effect can be used for magnetometry, but it requires a high time resolution of the detector. An easier characteristic to observe is the absorption profile averaged over the measurement time. We found that the averaged profile can be sensitive to both the strength and frequency of the test magnetic field, and this sensitivity is most pronounced for frequencies up to a hundred kHz and strengths up to several Gauss for the reference field of about one Gauss. Finally, Sec. IV provides a brief summary and outlook.
@@ II.2 Vector light beams
where $A_{0}=E_{0}/\omega$ and $a_{\varkappa m_{\gamma}}(\bm{k}_{\bot})$ is a weight function given by:
@@ II.3 Transition amplitudes
with $\bm{F}=\bm{I}+\bm{J}$, where $\bm{I}$ and $\bm{J}$ are the nuclear and electron angular momenta, respectively, $M$ is the projection of $\bm{F}$ on the quantization axis, $\alpha$ denotes all additional quantum numbers required to specify the state uniquely, $e$ is the elementary charge, and $c$ is the speed of light. Moreover, $q$ runs over all electrons in a target atom and $\bm{\alpha}_{q}$ denotes the vector of Dirac matrices for the $q$th particle Johnson (2007). In Eq. (8), we have introduced the impact parameter $\bm{b}=(b\cos\phi_{b},b\sin\phi_{b},0)$ to specify the position of the atom within the wave front Knyazev and Serbo (2018). In particular, $b=0$ corresponds to an atom located on the vortex line of a light beam. The impact parameter plays an important role in the analysis due to the complex spatial structure of the vector beam, see Fig. 3.
@@
where we have used the notation $\langle\alpha_{e}J_{e}||H_{\gamma}(pL)||\alpha_{g}J_{g}\rangle$ to denote the reduced matrix element for magnetic ($p=0$) and electric ($p=1$) transitions. Furthermore, the Euler angles as the arguments of the Wigner $D$-functions $d^{L}_{M,\lambda}(\theta_{k})$ and $D^{L}_{M_{e}-M_{g},M}(\pi,\pi/2,\pi-\phi_{B}(t))$ in Eq. (10) characterize the rotation from the atomic frame with the quantization axis along the magnetic field $\bm{B}_{\mathrm{total}}(t)$ to the photon frame with the quantization axis along the wave vector $\bm{k}$ Rose (1957). The time-dependent angle $\pi-\phi_{B}(t)$ refers to the oscillation of the polarization vector of the incident light at the position of the atom in its (time-dependent) reference frame. The transition amplitude (10) is valid for non-paraxial light beams as it was derived from the vector potential in the form (4) and (5). However, in order to explain the results of the present work we might restrict ourselves to the simpler case of paraxial light fields.